OpenStudy (anonymous):

48% of what # is 78???

OpenStudy (anonymous):

48% = 78 1% = 78/48 100% = 100 * 78/48

OpenStudy (anonymous):

so it would be 100 x 1 5/8?? which is 162 1/2 ?

OpenStudy (anonymous):

yeah i agree with that

OpenStudy (anonymous):

cool! :)